the area of a rectangle wall of a barn is 168 squate feet. its lenght is 10 feet longer than twice its width. find the lenght and width of the wall of the barn.
Answer by Cromlix(4381)

Hi there,
Make width = x
Length = 2x + 10
Area = Width * length (* = times)
168 = x(2x + 10)
168 = 2x^2 + 10x
2x^2 + 10x - 168 = 0
2(x^2 + 5x - 84) = 0
2(x + 12)(x - 7) = 0
x + 12 = 0
x = -12 (Disregard as -ve)
x - 7 = 0
x = 7
Width = 7 ft
Length = 24 ft
Hope this helps :-)